If only 8% of the energy supplied to a bulb is radiated as visible light then number of photons emitted per second by a 100 watt bulb are (Assume wavelength of visible light is 5.6×10^(-7)m )
1
Expert's answer
2021-09-08T16:13:47-0400

"λ=5.6 \times 10^{-7} \;m \\

h=6.625 \times 10^{-34} \;m^2gr/s"

Energy of one photone "= \frac{hc}{λ}"

"= \frac{6.625 \times 10^{-34} \times 3 \times 10^8}{5.6 \times 10^{-10}} \\

= 3.549 \times 10^{-19} \;J"

A 100 Watt bulb has to be supplied 100 J of energy.

Energy radiated by bulb per second as visible light "= \frac{100 \times 8}{100}= 8 \;J"

Visible light photons emitted per second as visible light "= \frac{8}{3.549 \times 10^{-19}}"

"=2.254 \times 10^{19 }"